OpenStudy (jkbo):

Help Please. What is the solution of the matrix equation? [31] [x] [3] [5 2] [y] = [10] A.(–6, 4) b.(6, –4) c.(4, 15) d.(–4, 15)

OpenStudy (anonymous):

The matrix equation : \[\begin{pmatrix}3&1\\5&2\end{pmatrix}\begin{pmatrix}x\\y\end{pmatrix}=\begin{pmatrix}3\\10\end{pmatrix}\] That means : \[\begin{cases}3x+y&=3~~~~~~~(1)\\5x+2y&=10~~~~~~(2)\end{cases}\] By multiplying the 1st equation by (1), we get : \[\begin{cases}6x+2y&=6~~~~~~~(1)\\5x+2y&=10~~~~~~(2)\end{cases}\] (1)-(2) : \[x=-4\] So : \[6(-4)+2y=6\] SO : \[-24+2y=6\] SO : \[2y=6+24\] SO : \[y=30/2=15\] So the answer is : D

OpenStudy (jkbo):

thank you

OpenStudy (jkbo):

and by multiplying the first equation by one does not give me 6 @Noura11

OpenStudy (agent0smith):

He meant to say multiply the first equation by 2.

OpenStudy (jkbo):

oo ok

OpenStudy (anonymous):

A mistype in the line 5 in my reply ! I am not a computer or a bike ! I am a human !

OpenStudy (jkbo):

sorry thank you though @Noura11

OpenStudy (anonymous):

I just was joking !

OpenStudy (agent0smith):

"I am not a computer or a bike ! I am a human !" Do bikes never make mistakes? :P